OpenStudy (anonymous):

steve traveled 200 miles at a certain speed. had he traveled 10mph faster, the trip would have taken 1 hour less. Find the speed of his vehicle.

ganeshie8 (ganeshie8):

\(\textbf{distance travelled = 200 miles} \\ \textbf{lets say his speed = x mph}\) now we can setup the equation : \[ \frac{200}{x}-1 = \frac{200}{x+10} \] x is your originial speed. solve to get the answer.. .

OpenStudy (anonymous):

how do I figure out what x is?

ganeshie8 (ganeshie8):

we got an equation with rational expressions... did they cover in ur syllabus... solving rational equations yet ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good.. so can you solve for x ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

don't fully understand yet.

ganeshie8 (ganeshie8):

ok no probs.. il walk you through this once...

OpenStudy (anonymous):

tks

ganeshie8 (ganeshie8):

\[\frac{200}{x}-1 = \frac{200}{x+10}\] \[200(x+10) -x(x+10) = 200x\] \[200x +2000 -x^2 -10x = 200x\] \[x^2+10x-2000 = 0\] \[(x+50)(x-40) = 0\] \[x=40\ mph\]

OpenStudy (anonymous):

Thank you, Appericate the help.

ganeshie8 (ganeshie8):

welcome :)